# Banach algebras associated to twisted étale groupoids: simplicity and pure infiniteness

### v2

@@ 4.1. Essential and exotic groupoid Banach algebras
Indeed, pick $\{f_{n}\}_{n=1}^{\infty}\subseteq\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ such that $\{i_{\mathcal{R}}(f_{n})\}_{n=1}^{\infty}$ converges to $f$ in $\|\cdot\|_{\mathcal{R}}$. Then $\{q_{\mathcal{G}}(f_{n})\}_{n=1}^{\infty}$ converges to $j_{\mathcal{R}}^{\operatorname{e}}(f)$ in the quotient norm of $\mathcal{D}(\mathcal{G},\mathcal{L})$. Hence also $g*q_{\mathcal{G}}(f_{n})\to g*j_{\mathcal{R}}^{\operatorname{e}}(f)$ by Lemma 4.11. Thus
@@
In view of Proposition 4.12, an exotic Banach algebra of $(\mathcal{G},\mathcal{L})$ is a Hausdorff completion $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$ of $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ equipped with a contractive linear map $j_{\mathcal{R}}^{\operatorname{e}}:F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})\to\mathcal{D}(\mathcal{G},\mathcal{L})$ which is identity on each of the spaces $C_{c}(U,\mathcal{L})$, $U\in\operatorname{Bis}(\mathcal{G})$. This map is injective if and only if $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$ is essential. By the same argument as in Corollary 3.9, for any exotic groupoid Banach algebra $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$ the quotient $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})/\ker j_{\mathcal{R}}^{\operatorname{e}}$ is naturally an essential groupoid Banach algebra of $(\mathcal{G},\mathcal{L})$. Every groupoid Banach algebra $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$ is exotic and we have $j_{\mathcal{R}}^{\operatorname{e}}=q_{\mathcal{G}}\circ j_{\mathcal{R}}$ and $\ker j_{\mathcal{R}}^{\operatorname{e}}=j_{\mathcal{R}}^{-1}(\mathfrak{M}_{0}(\mathcal{G},\mathcal{L}))$. In particular, if $\mathfrak{M}_{0}(\mathcal{G},\mathcal{L})=\{0\}$ then all reduced algebras are essential, as then $\ker j_{\mathcal{R}}^{\operatorname{e}}=\ker j_{\mathcal{R}}^{\operatorname{r}}$.
@@
Let $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$ be an exotic groupoid Banach algebra. For any representation $\psi:F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})\to B$ such that $\|q_{X}(f|_{X})\|\leq\|\psi(f)\|$ for $f\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$, we have $\ker\psi\subseteq\ker j_{\mathcal{R}}^{\operatorname{e}}$, and thus $\psi$ is necessarily injective on spaces $C_{c}(U,\mathcal{L})$, $U\in\operatorname{Bis}(\mathcal{G})$. If $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$ is an essential groupoid Banach algebra, then $\psi$ is necessarily injective on $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$.
@@
Consider the essential algebra $F^{p}_{\operatorname{e}}(\mathcal{G},\mathcal{L}):=F^{p}(\mathcal{G},\mathcal{L})/j_{p}^{-1}(\mathfrak{M}_{0}(\mathcal{G},\mathcal{L}))$ for $p\in[1,\infty]$. If $p=2$ then $F^{2}_{\operatorname{e}}(\mathcal{G},\mathcal{L})=C^{*}_{\operatorname{e}}(\mathcal{G},\mathcal{L})$ is the unique essential $C^{*}$-algebra. If $p\neq 2$, we do not know if $F^{\operatorname{e}}_{p}({\mathcal{G},\mathcal{L}})$ is an $L^{p}$-operator algebra, as such operator algebras in general are not closed under quotients. However, if $\mathcal{G}$ can be covered by a countable family of bisections (so for instance when $\mathcal{G}$ is $\sigma$-compact) or when $\mathcal{G}$ is topologically principal, then we have a natural injective representation of $F^{\operatorname{e}}_{p}({\mathcal{G},\mathcal{L}})$ on an $L^{p}$-space. Namely, by Lemma 4.4, $\mathcal{G}_{\operatorname{H}}$ is a full subgroupoid of $\mathcal{G}$, and so the regular representation $\Lambda_{p}:F^{p}(\mathcal{G},\mathcal{L})\to B(\ell^{p}(\mathcal{G},\mathcal{L}))$ compresses to an ‘essential representation’ $\widetilde{\Lambda}_{p}:F^{p}(\mathcal{G},\mathcal{L})\to B(\ell^{p}(\mathcal{G}_{\operatorname{H}},\mathcal{L}))$. By (3.8), we have
@@
As in the the proof of Proposition 3.15 we may assume that $p<\infty$. We identify $\widetilde{F}_{\operatorname{e}}^{p}({\mathcal{G},\mathcal{L}})$ with $\overline{\widetilde{\Lambda}_{p}(F^{p}(\mathcal{G},\mathcal{L}))}$, and we choose a Schauder basis $\{1_{\gamma}\}_{\gamma\in\mathcal{G}_{\operatorname{H}}}$ for $\ell^{p}(\mathcal{G}_{\operatorname{H}},\mathcal{L})$. By Lemma 4.4(2) the set $\mathcal{G}_{\operatorname{H}}$ is meager. Thus using (4.9) and recalling the notation from the proof of Proposition 3.15 for any $f\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ we get
@@
Hence we get a contractive linear map $j_{p}^{\operatorname{e}}:\widetilde{F}_{\operatorname{e}}^{p}({\mathcal{G},\mathcal{L}})\to\mathcal{D}(\mathcal{G},\mathcal{L})$ consistent with the quotient map $q_{\mathcal{G}}:\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\to\mathcal{D}(\mathcal{G},\mathcal{L})$. As in (3.8) one sees that for every $\gamma\in\mathcal{G}_{\operatorname{H}}$ and any $f\in\widetilde{F}_{\operatorname{e}}^{p}({\mathcal{G},\mathcal{L}})$ we have $\|f1_{\gamma}\|_{p}=\left(\sum_{d(\eta)=d(\gamma)}|j_{p}^{\operatorname{e}}(f)(\eta\gamma^{-1})|^{p}\right)^{1/p}$. Thus $j_{p}^{\operatorname{e}}$ is injective. ∎
@@
When $p=2$ the injective representation $\widetilde{\Lambda}_{p}$ is automatically isometric and so $\widetilde{F}_{\operatorname{e}}^{2}({\mathcal{G},\mathcal{L}})\cong F^{\operatorname{e}}_{2}({\mathcal{G},\mathcal{L}})\cong C^{*}_{\operatorname{e}}(\mathcal{G},\mathcal{L})$ is the unique essential $C^{*}$-algebra. This representation (for untwisted groupoids) appears also in [BM25, Proposition 3.31], and a similar representaion was constructed in [NS04, Proposition 1.12]. For $p\neq 2$ there is no reason for $\widetilde{\Lambda}_{p}$ to be isometric, unless $F^{\operatorname{r}}_{p}({\mathcal{G},\mathcal{L}})=F^{\operatorname{e}}_{p}({\mathcal{G},\mathcal{L}})$.
@@ 5. Intersection property and topological freeness
In this section we study conditions implying the following intersections properties for $A=C_{0}(X)$ and $B$ being an exotic groupoid Banach algebra, see [BK24, Definition 5.6] and [KM21, Definitions 5.5, 5.6]. Recall that the ideals we consider are closed and two-sided.
@@
Let $A\subseteq B$ be a Banach subalgebra of a Banach algebra $B$. We say that $A$ detects ideals in $B$, or that $A\subseteq B$ has the intersection property, if for every non-zero ideal $J$ in $B$ we have $J\cap A\neq\{0\}$. The inclusion has the generalized intersection property if there is a largest ideal $\operatorname{\mathcal{N}}$ in $B$ with $\operatorname{\mathcal{N}}\cap A=\{0\}$. In this case we call $\operatorname{\mathcal{N}}$ the hidden ideal and put $B_{\operatorname{e}}:=B/\operatorname{\mathcal{N}}$. We say $A\subseteq B$ is minimal if for every non-zero $a\in A$ we have $\overline{BaB}=B$.
@@
If $A\subseteq B$ has the generalized intersection property and $A$ is a $C^{*}$-algebra we have the embedding $A\subseteq B_{\operatorname{e}}$ is isometric (i.e. we have $A=\overline{A}$ in Lemma 5.2).
@@
Let $B$ be a Banach algebra of the form $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$ and such that $C_{0}(X)$ embeds isometrically into $B$. For any ideal $I$ in $C_{0}(X)$ the following are equivalent:
@@
    $I$ is symmetric, i.e. $\overline{IB}=\overline{BI}$.
In particular, the inclusion $C_{0}(X)\subseteq B$ is minimal if and only if $\mathcal{G}$ is minimal.
@@
We have $I=C_{0}(U)$ for an open set $U\subseteq X$. For $V\in\operatorname{Bis}(\mathcal{G})$ we denote by $B_{V}$ the image of $C_{0}(V,\mathcal{L})$ in $B=F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$.
(1$\Rightarrow$(2). Assume $I=C_{0}(X)\cap J$ for an ideal $J$ in $B$. For every $V\in\operatorname{Bis}(\mathcal{G})$ we have $C_{0}(VUV^{*})=\overline{B_{V}C_{0}(U)B_{V^{*}}}\subseteq J\cap C_{0}(X)=C_{0}(U)$. Hence $VUV^{*}\subseteq U$ for every $V\in\operatorname{Bis}(\mathcal{G})$, which implies $U$ is $\mathcal{G}$-invariant.
(2$\Rightarrow$(3). If (2) holds then $VU=UV$ for all $V\in\operatorname{Bis}(\mathcal{G})$, which implies that $IB_{V}=B_{V}I$ for $V\in\operatorname{Bis}(\mathcal{G})$. Hence $\overline{IB}=\overline{\sum_{V\in\operatorname{Bis}(\mathcal{G})}IB_{V}}=\overline{\sum_{V\in\operatorname{Bis}(\mathcal{G})}B_{V}I}=\overline{BI}$.
(3$\Rightarrow$(1). If (3) holds, then $J=\overline{IB}=\overline{BI}$ is an ideal in $B$ (generated by $I$) and an approximate unit $\{\mu_{i}\}_{i}$ in $I$ is an approximate unit in $J$. Using this, one gets $I=J\cap C_{0}(X)$.
Now assume $\mathcal{G}$ is minimal and let $a\subseteq C_{0}(X)$ be non-zero. By the equivalence (1$\Leftrightarrow$(2), $\overline{BaB}\cap C_{0}(X)=C_{0}(U)$ where $U$ is an open $\mathcal{G}$-invariant set. Since $U$ contains the open support of $a$ it is non-empty. Hence $U=X$ by $\mathcal{G}$-minimality, and therefore $C_{0}(X)\subseteq\overline{BaB}$. Since $C_{0}(X)$ contains an approximate unit for $B$ this implies that $\overline{BaB}=B$. Thus $C_{0}(X)\subseteq B$ is minimal. For the converse assume $\mathcal{G}$ is not minimal. Let $U\subseteq X$ be a non-trivial open $\mathcal{G}$-invariant set and put $I=C_{0}(U)$. By the equivalence (2$\Leftrightarrow$(3), $\overline{BIB}=\overline{IB}$ and so any approximate unit in $I$ is also an approximate unit in $\overline{BIB}$. This implies that $\overline{BIB}\cap C_{0}(X)=I\neq C_{0}(X)$. Hence for any non-zero $a\in I$ we get $\overline{BaB}\neq B$. Thus $C_{0}(X)\subseteq B$ is not minimal. ∎
@@
The groupoid $\mathcal{G}$ is topologically free if and only if for every bisection $V\subseteq\mathcal{G}\setminus X$, the set $\{x\in X:\mathcal{G}(x)\cap V\neq\emptyset\}=\{x\in d(V):h_{V}(x)=x\}$ has empty interior in $X$ ($h_{V}=r\circ d|_{V}^{-1}$ is the partial homeomorphism induced by $V$). The relationship between this and other popular non-triviality conditions is presented on the following diagram:
@@
Effectivness of $\mathcal{G}$ means that for any open $V\subseteq\mathcal{G}$ with $r|_{V}=d|_{V}$ we have $V\subseteq X$. It implies topological freeness and the converse is true if $\mathcal{G}$ is Hausdorff ($X$ is closed in $\mathcal{G}$). If $\mathcal{G}$ is topologically principal, that is if $\{x\in X:\mathcal{G}(x)\neq\emptyset\}$ has empty interior, then $\mathcal{G}$ is topologically free and the converse is true if $\mathcal{G}$ is covered by a countable family of bisections (so e.g. when $\mathcal{G}$ is second countable or more generally $\sigma$-compact). In general, all these conditions are different and topological freeness is the weakest. See [KM21, Section 2.4] for more details.
The following lemma is the technical heart of our simplicity and pure infiniteness criteria. For these results we will need an $S$-grading, that we have been able to ignore until now. In what follows we suppress writing the convolution sign ‘*’ when multiplying elements of $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$.
@@
Assume $\mathcal{G}$ is topologically free and let $S\subseteq\operatorname{Bis}(\mathcal{G})$ be a unital inverse semigroup covering $\mathcal{G}$. For any $f\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ and any $\varepsilon>0$ there is a contractive element $b\in C_{c}(X)$ and a bump function $a\in C_{c}(X)$ such that $bf|_{X}:=(bf)|_{X}=b(f|_{X})$ is in $C_{0}(X)^{+}$ and
@@
has empty interior. Thus there is $x_{0}\in W\setminus R$. For each $V\in F$ we claim there is a bump function $a_{V}\in C_{c}(W)$ equal to $1$ in a neighbourhood of $x_{0}$ and
(5.12) $|a_{V}(r(\gamma))\cdot f_{V}(\gamma)\cdot a_{V}(d(\gamma))|\leq\varepsilon/|F|\qquad\gamma\in V\setminus\overline{X}.$
@@
(1) $x_{0}\in X\setminus r\left(\overline{\operatorname{supp}(f_{V})}\right)$. Then $f_{V}(r|_{V}^{-1}(x_{0}))=0$ so there is a neighbourhood $U\subseteq W$ of $x_{0}$ where $|f_{V}\circ(r|_{V})^{-1}|\leq\varepsilon/|F|$. Taking any norm one $a_{V}\in C_{0}(U)$ we get $|a_{V}(r(\gamma))\cdot f_{V}(\gamma)|\leq\varepsilon/|F|$ for $\gamma\in V\setminus\overline{X}$, and so (5.12) is satisfied.
(2) $x_{0}\in r\left(\overline{\operatorname{supp}(f_{V})}\cap X\right)\subseteq r(V\cap X)=V\cap X$. Thus $x_{0}\in V\cap W$ and for any $a_{V}\in C_{c}(V\cap W)$ we get $a_{V}(r(\gamma))\cdot f_{V}(\gamma)=0$ for $\gamma\in V\setminus\overline{X}$ because $d(V\setminus\overline{X})\cap V=\emptyset$.
(3) $x_{0}\in r\left(\overline{\operatorname{supp}(f_{V})}\setminus\overline{X}\right)\subseteq r(V\setminus\overline{X})$. Since $x_{0}\not\in R$ there is a neighborhood $U\subseteq r(V\setminus\overline{X})\cap W$ of $x_{0}$ such that $h_{V}^{-1}(U)\cap U=\emptyset$. Then for any $a_{V}\in C_{0}(U)$ we have $a_{V}(r(\gamma))a_{V}(d(\gamma))=0$ for $\gamma\in V\setminus\overline{X}$, and so (5.12) holds.
This proves our claim. Since $W\cap r\left(\partial X\cap\overline{\operatorname{supp}(f_{V})}\right)=\emptyset$ we have $(a_{V}f_{V})|_{\partial X}=0$. Thus $(a_{V}f_{V})|_{X}=a_{V}(f_{V}|_{X})\in C_{0}(X)$, $(a_{V}f_{V})|_{V\setminus\overline{X}}\in C_{0}(V\setminus\overline{X},\mathcal{L})\subseteq\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ and $a_{V}f_{V}=(a_{V}f_{V})|_{X}+(a_{V}f_{V})|_{V\setminus\overline{X}}$. Therefore putting $a:=\prod_{V\in F}a_{V}\in C_{c}(W)$ we get a bump function equal to $1$ in a neighborhood of $x_{0}$ such that
@@
Since $\psi|_{C_{0}(X)}$ is injective, $\psi|_{C_{0}(X)}$ is isometric by minimality of the supremum norm, cf. [BKM23, Proposition 2.10]. Let $f\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ and take $\epsilon>0$. For $a,b\in C_{0}(X)$ as in Lemma 5.8, using the inequalities there, we get

### v3

@@ 4.2. Essential and exotic groupoid Banach algebras
Indeed, pick a sequence $(f_{n})_{n=1}^{\infty}$ in $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ such that $(i_{\mathcal{R}}(f_{n}))_{n=1}^{\infty}$ converges to $f$ in $\|\cdot\|_{\mathcal{R}}$. Then $(q_{\mathcal{G}}(f_{n}))_{n=1}^{\infty}$ converges to $j_{\mathcal{R}}^{\operatorname{e}}(f)$ in the quotient norm of $\mathcal{D}(\mathcal{G},\mathcal{L})$. Hence also $g*q_{\mathcal{G}}(f_{n})\to g*j_{\mathcal{R}}^{\operatorname{e}}(f)$ by Lemma 4.11. Thus
@@
In view of Proposition 4.12, an exotic Banach algebra of $(\mathcal{G},\mathcal{L})$ is a Hausdorff completion $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$ of $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ equipped with a contractive linear map $j_{\mathcal{R}}^{\operatorname{e}}:F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})\to\mathcal{D}(\mathcal{G},\mathcal{L})$ which is the identity on each of the spaces $C_{c}(U,\mathcal{L})$, $U\in\operatorname{Bis}(\mathcal{G})$. This map is injective if and only if $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$ is essential. By the same argument as in Corollary 3.9, for any exotic groupoid Banach algebra $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$, the quotient $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})/\ker j_{\mathcal{R}}^{\operatorname{e}}$ is naturally an essential groupoid Banach algebra of $(\mathcal{G},\mathcal{L})$. Every groupoid Banach algebra $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$ is exotic and we have $j_{\mathcal{R}}^{\operatorname{e}}=q_{\mathcal{G}}\circ j_{\mathcal{R}}$ and $\ker j_{\mathcal{R}}^{\operatorname{e}}=j_{\mathcal{R}}^{-1}(\mathfrak{M}_{0}(\mathcal{G},\mathcal{L}))$. In particular, if $\mathfrak{M}_{0}(\mathcal{G},\mathcal{L})=\{0\}$, then all reduced algebras are essential, as then $\ker j_{\mathcal{R}}^{\operatorname{e}}=\ker j_{\mathcal{R}}^{\operatorname{r}}$.
@@
Let $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$ be an exotic groupoid Banach algebra. For any representation $\psi:F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})\to B$ such that $\|q_{X}(f|_{X})\|\leq\|\psi(f)\|$ for $f\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$, we have $\ker\psi\subseteq\ker j_{\mathcal{R}}^{\operatorname{e}}$, and thus $\psi$ is necessarily injective on the spaces $C_{c}(U,\mathcal{L})$, $U\in\operatorname{Bis}(\mathcal{G})$. If $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$ is an essential groupoid Banach algebra, then $\psi$ is necessarily injective on $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$.
@@
Consider the essential algebra $F^{p}_{\operatorname{e}}(\mathcal{G},\mathcal{L}):=F^{p}(\mathcal{G},\mathcal{L})/j_{p}^{-1}(\mathfrak{M}_{0}(\mathcal{G},\mathcal{L}))$ for $p\in[1,\infty]$. If $p=2$, then $F^{2}_{\operatorname{e}}(\mathcal{G},\mathcal{L})=C^{*}_{\operatorname{e}}(\mathcal{G},\mathcal{L})$ is the unique essential $C^{*}$-algebra. If $p\neq 2$, we do not know if $F^{\operatorname{e}}_{p}({\mathcal{G},\mathcal{L}})$ is an $L^{p}$-operator algebra, as such operator algebras in general are not closed under quotients. However, if $\mathcal{G}$ can be covered by a countable family of bisections (so for instance when $\mathcal{G}$ is $\sigma$-compact) or when $\mathcal{G}$ is topologically principal, then we have a natural injective representation of $F^{\operatorname{e}}_{p}({\mathcal{G},\mathcal{L}})$ on an $L^{p}$-space. Namely, by Lemma 4.4, $\mathcal{G}_{\operatorname{H}}$ is a full subgroupoid of $\mathcal{G}$, and so the regular representation $\Lambda_{p}:F^{p}(\mathcal{G},\mathcal{L})\to B(\ell^{p}(\mathcal{G},\mathcal{L}))$ compresses to an ‘essential representation’ $\widetilde{\Lambda}_{p}:F^{p}(\mathcal{G},\mathcal{L})\to B(\ell^{p}(\mathcal{G}_{\operatorname{H}},\mathcal{L}))$. By (3.7), we have
@@
As in the the proof of Proposition 3.15, we may assume that $p<\infty$. We identify $\widetilde{F}_{\operatorname{e}}^{p}({\mathcal{G},\mathcal{L}})$ with $\overline{\widetilde{\Lambda}_{p}(F^{p}(\mathcal{G},\mathcal{L}))}$, and we choose a Schauder basis $\{1_{\gamma}\}_{\gamma\in\mathcal{G}_{\operatorname{H}}}$ for $\ell^{p}(\mathcal{G}_{\operatorname{H}},\mathcal{L})$. By Lemma 4.4(2), the set $\mathcal{G}_{\operatorname{H}}$ is meager. Thus, using (4.8) and recalling the notation from the proof of Proposition 3.15, for any $f\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ we get
@@
Hence we get a contractive linear map $j_{p}^{\operatorname{e}}:\widetilde{F}_{\operatorname{e}}^{p}({\mathcal{G},\mathcal{L}})\to\mathcal{D}(\mathcal{G},\mathcal{L})$ consistent with the quotient map $q_{\mathcal{G}}:\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\to\mathcal{D}(\mathcal{G},\mathcal{L})$. As in (3.7) one sees that for every $\gamma\in\mathcal{G}_{\operatorname{H}}$ and any $f\in\widetilde{F}_{\operatorname{e}}^{p}({\mathcal{G},\mathcal{L}})$, we have $\|f1_{\gamma}\|_{p}=\left(\sum_{d(\eta)=d(\gamma)}|j_{p}^{\operatorname{e}}(f)(\eta\gamma^{-1})|^{p}\right)^{1/p}$. Thus $j_{p}^{\operatorname{e}}$ is injective. ∎
@@
When $p=2$, the injective representation $\widetilde{\Lambda}_{p}$ is automatically isometric and so $\widetilde{F}_{\operatorname{e}}^{2}({\mathcal{G},\mathcal{L}})\cong F^{\operatorname{e}}_{2}({\mathcal{G},\mathcal{L}})\cong C^{*}_{\operatorname{e}}(\mathcal{G},\mathcal{L})$ is the unique essential $C^{*}$-algebra. This representation (for untwisted groupoids) also appears in [BM25, Proposition 3.31], and a similar representation was constructed in [NS04, Proposition 1.12]. For $p\neq 2$, there is no reason for $\widetilde{\Lambda}_{p}$ to be isometric, unless $F^{\operatorname{r}}_{p}({\mathcal{G},\mathcal{L}})=F^{\operatorname{e}}_{p}({\mathcal{G},\mathcal{L}})$.
@@ 5. Intersection property and topological freeness
In this section, we study conditions implying the following intersection properties for $A=C_{0}(X)$ and $B$ being an exotic groupoid Banach algebra, see [BK24, Definition 5.6] and [KM21, Definitions 5.5, 5.6].
5.1. Intersection property and topological freeness
Recall that the ideals we consider are closed and two-sided.
@@ 5.1. Intersection property and topological freeness
Let $A\subseteq B$ be a Banach subalgebra of a Banach algebra $B$. We say that $A$ detects ideals in $B$, or that $A\subseteq B$ has the intersection property, if for every non-zero ideal $J$ in $B$, we have $J\cap A\neq\{0\}$. The inclusion has the generalized intersection property if there is a largest ideal $\operatorname{\mathcal{N}}$ in $B$ with $\operatorname{\mathcal{N}}\cap A=\{0\}$. In this case we call $\operatorname{\mathcal{N}}$ the hidden ideal and put $B_{\operatorname{e}}:=B/\operatorname{\mathcal{N}}$. We say $A\subseteq B$ is minimal if for every non-zero $a\in A$ we have $\overline{BaB}=B$.
@@
If $A\subseteq B$ has the generalized intersection property and $A$ is a $C^{*}$-algebra, then the embedding $A\subseteq B_{\operatorname{e}}$ is isometric (that is, we have $A=\overline{A}$ in Lemma 5.2).
@@
Let $B$ be a Banach algebra of the form $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$ and such that $C_{0}(X)$ embeds isometrically into $B$. For any ideal $I$ in $C_{0}(X)$, the following are equivalent:
@@
    II is symmetric, that is, $\overline{IB}=\overline{BI}$.
Moreover, the inclusion $C_{0}(X)\subseteq B$ is minimal if and only if $\mathcal{G}$ is minimal.
@@
We have $I=C_{0}(U)$ for an open set $U\subseteq X$. For $V\in\operatorname{Bis}(\mathcal{G})$, we denote the image of $C_{0}(V,\mathcal{L})$ in $B=F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$ by $B_{V}$.
[…]
@@
The groupoid $\mathcal{G}$ is topologically free if and only if for every bisection $V\subseteq\mathcal{G}\setminus X$, the set $\{x\in X:\mathcal{G}(x)\cap V\neq\emptyset\}=\{x\in d(V):h_{V}(x)=x\}$ has empty interior in $X$ ($h_{V}=r\circ d|_{V}^{-1}$ is the partial homeomorphism induced by $V$). The relationship between this and other popular non-triviality conditions is presented in the following diagram:
@@
Effectiveness of $\mathcal{G}$ means that for any open $V\subseteq\mathcal{G}$ with $r|_{V}=d|_{V}$ we have $V\subseteq X$. It implies topological freeness and the converse is true if $\mathcal{G}$ is Hausdorff ($X$ is closed in $\mathcal{G}$). If $\mathcal{G}$ is topologically principal, that is if $\{x\in X:\mathcal{G}(x)\neq\{x\}\}$ has empty interior, then $\mathcal{G}$ is topologically free and the converse is true if $\mathcal{G}$ is covered by a countable family of bisections (so, for example, when $\mathcal{G}$ is second countable or more generally $\sigma$-compact). In general, all these conditions are different and topological freeness is the weakest. See [KM21, Section 2.4] for more details.
The following lemma is the technical heart of our simplicity and pure infiniteness criteria. For these results, we will need an $S$-grading, that we have been able to ignore until now. In what follows, we suppress writing the convolution sign ‘*’ when multiplying elements of $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$.
@@
Assume $\mathcal{G}$ is topologically free and let $S\subseteq\operatorname{Bis}(\mathcal{G})$ be a unital inverse semigroup covering $\mathcal{G}$. For any $f\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ and any $\varepsilon>0$, there is a contractive element $b\in C_{c}(X)$ and a bump function $a\in C_{c}(X)$ such that $bf|_{X}:=(bf)|_{X}=b(f|_{X})$ is in $C_{0}(X)^{+}$ and
@@
has empty interior. Thus there is $x_{0}\in W\setminus R$. For each $V\in F$, we claim there is a bump function $a_{V}\in C_{c}(W)$ equal to $1$ in a neighbourhood of $x_{0}$ and
(5.11) $|a_{V}(r(\gamma))\cdot f_{V}(\gamma)\cdot a_{V}(d(\gamma))|\leq\varepsilon/|F|,\qquad\gamma\in V\setminus\overline{X}.$
@@
(1) $x_{0}\in X\setminus r\left(\overline{\operatorname{supp}(f_{V})}\right)$. Then $f_{V}(r|_{V}^{-1}(x_{0}))=0$ so there is a neighbourhood $U\subseteq W$ of $x_{0}$ where $|f_{V}\circ(r|_{V})^{-1}|\leq\varepsilon/|F|$. Taking any norm one $a_{V}\in C_{0}(U)$, we get $|a_{V}(r(\gamma))\cdot f_{V}(\gamma)|\leq\varepsilon/|F|$ for $\gamma\in V\setminus\overline{X}$, and so (5.11) is satisfied.
(2) $x_{0}\in r\left(\overline{\operatorname{supp}(f_{V})}\cap X\right)\subseteq r(V\cap X)=V\cap X$. Thus $x_{0}\in V\cap W$ and, for any $a_{V}\in C_{c}(V\cap W)$, we get $a_{V}(r(\gamma))\cdot f_{V}(\gamma)=0$ for $\gamma\in V\setminus\overline{X}$ because $d(V\setminus\overline{X})\cap V=\emptyset$.
(3) $x_{0}\in r\left(\overline{\operatorname{supp}(f_{V})}\setminus\overline{X}\right)\subseteq r(V\setminus\overline{X})$. Since $x_{0}\not\in R$, there is a neighborhood $U\subseteq r(V\setminus\overline{X})\cap W$ of $x_{0}$ such that $h_{V}^{-1}(U)\cap U=\emptyset$. Then, for any $a_{V}\in C_{0}(U)$, we have $a_{V}(r(\gamma))a_{V}(d(\gamma))=0$ for $\gamma\in V\setminus\overline{X}$, and so (5.11) holds.
This proves our claim. Since $W\cap r\left(\partial X\cap\overline{\operatorname{supp}(f_{V})}\right)=\emptyset$, we have $(a_{V}f_{V})|_{\partial X}=0$. Thus $(a_{V}f_{V})|_{X}=a_{V}(f_{V}|_{X})\in C_{0}(X)$, $(a_{V}f_{V})|_{V\setminus\overline{X}}\in C_{0}(V\setminus\overline{X},\mathcal{L})\subseteq\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ and $a_{V}f_{V}=(a_{V}f_{V})|_{X}+(a_{V}f_{V})|_{V\setminus\overline{X}}$. Therefore putting $a:=\prod_{V\in F}a_{V}\in C_{c}(W)$, we get a bump function equal to $1$ in a neighborhood of $x_{0}$ such that
@@
Since $\psi|_{C_{0}(X)}$ is injective, $\psi|_{C_{0}(X)}$ is isometric by minimality of the supremum norm, cf. [BKM25, Proposition 2.10]. Let $f\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ and take $\epsilon>0$. For $a,b\in C_{0}(X)$ as in Lemma 5.8, using the inequalities there, we get